OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table? x~ -2 , 0 , 4 f(x)~ 1 , 5 , -59

OpenStudy (anonymous):

the general form of a parabola is f(x) = ax^2+bx+c now we have to satisfy the conditions : f(-2) = 1 f(0) = 5 f(4) = -59 using this we will find a,b,c.

OpenStudy (anonymous):

so it will be: a(-2)^2+b(-2)+c ?

OpenStudy (anonymous):

f(-2) = 1 means a(-2)^2+b(-2)+c = 1

OpenStudy (anonymous):

we will have 3 equations so we could find a,b and c.

OpenStudy (anonymous):

and the 2nd will be : a(0)^2+b(0)+c=5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

write it after you simplify ..

OpenStudy (anonymous):

and the 3rd will be : a(4)^2+b(4)+c=-59

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so write the 3 simplified equations

OpenStudy (anonymous):

so the 1st one simplified is a*4+b*-2+c=1 ?

OpenStudy (anonymous):

yes which is 4a-2b+c=1

OpenStudy (anonymous):

now what are the other 2 ?

OpenStudy (anonymous):

the 2nd would be : 0a+0b+c=5 ?

OpenStudy (anonymous):

which means ? what is 0a ? 0b ?

OpenStudy (anonymous):

what ?

OpenStudy (anonymous):

0*a is 0a which is 0 .. since it's 0 times something .. so 0a +0b+c=5 is like c = 5 agree?

OpenStudy (anonymous):

agree (:

OpenStudy (anonymous):

4a-2b+c=1 c = 5 what is the third ?

OpenStudy (anonymous):

& the third would be: 16a+4b+c=-59

OpenStudy (anonymous):

good so we have system of equations : 4a-2b+c=1 c = 5 16a+4b+c=-59

OpenStudy (anonymous):

we know from the second equation that c = 5.. can you find a and b ?

OpenStudy (anonymous):

do i plug 5 for c in the 1st & 3rd equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then you left with 2 equations and 2 unknowns

OpenStudy (anonymous):

so : 4a-2b+5=1 ?

OpenStudy (anonymous):

this is the first

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you solve 2 equations with 2 unknowns ? find a and b ..

OpenStudy (anonymous):

I dont know how to find a & b ?

OpenStudy (anonymous):

@andreascott "do i plug 5 for c in the 1st & 3rd equation?" Yes, please do ! then show your results.

OpenStudy (anonymous):

so : 4a-2b+5=1 ?

OpenStudy (anonymous):

what about the third ?

OpenStudy (anonymous):

16a+4b+5=-59 ?

OpenStudy (anonymous):

4a-2b+5=1 --> 4a - 2b = ...?

OpenStudy (anonymous):

im confused now ?

OpenStudy (anonymous):

we want in the left side of the equations only a,b so we want to get rid of the 5 in the left side what the equations will look like ?

OpenStudy (anonymous):

4a-2b ?

OpenStudy (anonymous):

= ?

OpenStudy (anonymous):

4a-2b = 5 ??

OpenStudy (anonymous):

4a-2b+5=1 subtract 5 from both sides (in order to have only a and b in the left side) 4a-2b +5 -5 = 1-5 4a-2b = -4

OpenStudy (anonymous):

can you do the same for the other equation ?

OpenStudy (anonymous):

16a+4b+5=-59 ->16a+4b = ...?

OpenStudy (anonymous):

16a+4b+5=-59 ? 16a+4b+5 - 5 = 20 -5 ? 16a+4b = 15 ?

OpenStudy (anonymous):

how -59 turned into 20 ?

OpenStudy (anonymous):

i thought you did 4-2 = 1 so i did 16+4 =20 ?

OpenStudy (anonymous):

first, 4-2 = 2 second you have to work with the original equation which is (ill show work on the first one) 4a-2b+5=1 now i subtract 5 from both sides of the equation (the 1 was there because this is the original equation) 4a-2b +5-5= 1-5 4a-2b = -4

OpenStudy (anonymous):

4a-2b +5 -5 = 1 -5 4a-2b = -4

OpenStudy (anonymous):

i havent changed any thing .. just moved the 5 into the right side

OpenStudy (anonymous):

ohhhhhhhh , i feel reallly slow . excuse me . lol im not thinking right

OpenStudy (anonymous):

it's fine

OpenStudy (anonymous):

im sure you done those things before

OpenStudy (anonymous):

can you do the same for the second equation now ?

OpenStudy (anonymous):

I have a head -ache -______-

OpenStudy (anonymous):

16a+4b+5 -5 = -59 -5 16a+4b = -64 ?

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

so now we have 4a-2b = -4 16a+4b = -64 in order to solve for a or for b we can use the elimination or substitution method have you heard about it ?

OpenStudy (anonymous):

heard about it , but cant remember how to start it off ?

OpenStudy (anonymous):

for example .. if we want to use the elimination method and we want to eliminate a and solve for b first we may multiply both sides of the first equation by 4 (so the a term will be 16a) and then we will have : 4(4a-2b) = -4 *4 16a - 8b = -16 so we have 16a - 8b = -16 16a+4b = -64 and then we can subtract the second from the first can you show me how you do this ?

OpenStudy (anonymous):

plug -64 into 4a-2b = -4 ?

OpenStudy (anonymous):

we have to subtract the second equation from the first so we subtract the right side of the second equation from the right side of the first and subtract the left side of the second equation from the left side of the first

OpenStudy (anonymous):

i dont understand ?

OpenStudy (anonymous):

16a - 8b = -16 16a+4b = -64 -> 16a - 8b -(16a+4b) = -16 - (-64)

OpenStudy (anonymous):

i kinda see . . .

OpenStudy (anonymous):

now simplify

OpenStudy (anonymous):

a -4b = 48 ?

OpenStudy (anonymous):

no show me how you did it

OpenStudy (anonymous):

i did 16a -16a & got 0 . then i did -8b + 4b & got -4 . then i did -16 - ( 64) & got 48 .

OpenStudy (anonymous):

so why i see a in your answer

OpenStudy (anonymous):

and it is -8b-4b

OpenStudy (anonymous):

ohhh , there is no a . my fault . so its -4b=48 ?

OpenStudy (anonymous):

and it is -8b-4b

OpenStudy (anonymous):

& when u simplify that its -12b ?

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

so what is the equation after simplification

OpenStudy (anonymous):

-12b=48

OpenStudy (anonymous):

what is b then?

OpenStudy (anonymous):

i knew that , im just thinking tooo hard .

OpenStudy (anonymous):

b=-4 ?

OpenStudy (anonymous):

good.. so now in order to find a we plug b in one of the two equations we had .

OpenStudy (anonymous):

4a-2(-4) = -4

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes so what is a ?

OpenStudy (anonymous):

8 or 2 ?

OpenStudy (anonymous):

there is only one answer show me

OpenStudy (anonymous):

4a+8=-4 ?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

so what is a

OpenStudy (anonymous):

idk ? 4 ?

OpenStudy (anonymous):

4a+8=-4 isolate a .. subtract 8 from both sides .. its one equation with one unknown you know how to solve for a!

OpenStudy (anonymous):

8-4a +8 = -4-8 ?

OpenStudy (anonymous):

4a+8=-4 subtract 8 from both sides 4a +8-8 = -4-8 4a = -12 divide both sides by 4 a = -3

OpenStudy (anonymous):

you know how to do it .. Im sure.. i guess you are tired.

OpenStudy (anonymous):

so we found a,b and c look at the first thing i wrote in this post and tell me what is the parabola

OpenStudy (anonymous):

f(x) = -3x^2+-4x+5 ?

OpenStudy (anonymous):

+-4x is like -4x

OpenStudy (anonymous):

f(x) = -3x^2-4x+5 ?

OpenStudy (anonymous):

yes.. a good thing to do is to check ourselves and see if we plug in the data we get the correct given data